OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? A) 1/5 B)-1/5 C)5 D)-5

OpenStudy (anonymous):

-1/5

OpenStudy (anonymous):

so B :)

OpenStudy (bahrom7893):

M = (y2-y1)/(x2-x1) = (4-5)/(6-1) = -1/5

OpenStudy (anonymous):

The slope is the defined as the rate of change between two things. This is defined as\[m = (\Delta y/\Delta x)\] where m is the slope.and \[\Delta y = y2-y1\] which represents the change in y. The same applies to \[\Delta x \].